OpenStudy (anonymous):

A tennis ball is thrown vertically upward with an initial velocity of +9.3 m/s. What will the ball’s velocity be when it returns to its starting point? The acceleration of gravity is 9.81 m/s^2 How long will the ball take to reach its starting point?

OpenStudy (roadjester):

It should have the same velocity. Projectile motion is like a parabola.

OpenStudy (anonymous):

Since the ball is launched vertically against gravity force at 9.3 m/s, then it will also fall back down from the peak and have the same velocity of 9.3 m/s at ground start point equations: y = Voy * t - 1/2 g * t^2 and Vy = Voy - gt Vy = velocity of object in y direction Voy = initial vertical velocity t = time in seconds y = distance in vertical meters g = 9.81 m/s^2 Solving for y, the object goes 4.4 meters high solving for t: t = 0.95 seconds for trajectory time